# Test-Time 3D Occupancy Prediction

### v3

@@ Abstract
Self-supervised 3D occupancy prediction offers a promising solution for understanding complex driving scenes without requiring costly 3D annotations. However, training dense occupancy decoders to capture fine-grained geometry and semantics can demand hundreds of GPU hours, and once trained, such models struggle to adapt to varying voxel resolutions or novel object categories without extensive retraining. To overcome these limitations, we propose a practical and flexible test-time occupancy prediction framework termed TT-Occ. Our method incrementally constructs, optimizes and voxelizes time-aware 3D Gaussians from raw sensor streams by integrating vision foundation models (VFMs) at runtime. The flexible nature of 3D Gaussians allows voxelization at arbitrary user-specified resolutions, while the generalization ability of VFMs enables accurate perception and open-vocabulary recognition, without any network training or fine-tuning. Specifically, TT-Occ operates in a “lift-track-voxelize” symphony: We first “lift” the geometry and semantics of surrounding-view extracted from VFMs to instantiate Gaussians at 3D space; Next, we “track” dynamic Gaussians while accumulating static ones to complete the scene and enforce temporal consistency; Finally, we voxelize the optimized Gaussians to generate occupancy prediction. Optionally, inherent noise in VFM predictions and tracking is mitigated by periodically smoothing neighboring Gaussians during optimization. To validate the generality and effectiveness of our framework, we offer two variants: one LiDAR-based and one vision-centric, and conduct extensive experiments on Occ3D and nuCraft benchmarks with varying voxel resolutions.
@@ 1 Introduction
3D Occupancy prediction seeks to accurately identify regions within an environment that are occupied by objects of particular classes and those that remain free. This capability is crucial to enable collision-free trajectory planning and reliable navigation in autonomous driving systems [planning_ad1, planning_ad2] and embodied agents [planning_agent, planning_agent2, navi]. Existing solutions [TPVFormer, OccFormer, RenderOcc, OccFlowNet, SurroundOcc, GaussianFormer, Occ3D], however, heavily rely on dense 3D annotations obtained through labor-intensive manual labeling of dynamic driving scenes spanning up to 80 meters per frame. To mitigate this cost, recent studies have resorted to self-supervised alternatives [SimpleOcc, OccNeRF, SelfOcc, DistillNeRF, GaussianOcc, GaussianTR, VEON, LangOcc, zhou2025autoocc]. While reducing labeling costs, they still require substantial computational overhead. For instance, training SelfOcc [SelfOcc] on Occ3D-nuScenes [Occ3D] at a voxel resolution of 0.4m requires approximately 2 days on eight GPUs. Furthermore, once trained, adapting to finer resolution (e.g., 0.2m of nuCraft [nuCraft] dataset) or novel object classes (e.g., beyond the 17 predefined classes of Occ3D-nuScenes [Occ3D]) may necessitate extensive retraining.
@@
Recent vision foundation models (VFMs), however, change this landscape. Models such as VGGT [vggt] and MapAnything [keetha2025mapanything] provide reliable multiview geometry, and REX-Omni [rexomni] enables open-vocabulary semantic reasoning, both directly at test time. As these cues become increasingly accessible without task-specific training, it is natural to revisit a long-standing assumption of the field: if such information no longer needs to be learned by a network, is training an occupancy model still necessary? To answer this question, we explore a test-time occupancy estimation method termed TT-Occ, which progressively constructs, optimizes and voxelizes time-aware 3D Gaussians from raw sensor streams by integrating VFMs. We introduce two variants, TT-OccCamera and TT-OccLiDAR, which differ in the sensor modality used to initialize the Gaussian primitives, respectively. Both variants eliminate costly pretraining and allow flexible adaptation to arbitrary voxel resolutions and user-specified semantic queries. Unlike previous NeRF- [emernerf, unisim] and 3DGS-based [DrivingGaussian, DriveStudio] reconstruction methods that perform offline per-scene modeling assisted by external GT priors (e.g., HD maps or bounding boxes), TT-Occ generates occupancy representations in an online fashion, relying solely on raw sensor streams and generally trained VFMs to instantiate Gaussians capturing object geometry and semantics in unbounded outdoor scenes.
@@
Our framework follows a simple yet effective “lift-track-voxelize” symphony: (1) Lift: at each test time step, we first “lift” geometry and semantic information of surrounding views extracted via VFMs into time-aware 3D Gaussians on the fly. The generated Gaussians can also be splatted back onto the image plane through differentiable rasterization for parameter optimization [3DGS]. (2) Track: next, we “track” dynamic Gaussians and accumulate static ones using estimated motion flow. This motion compensates for partial object visibility and prevents trailing artifacts while maintaining long-term temporal coherence. (3) Voxelize: at any given timestamp, the generated 3D Gaussians can be voxelized onto discrete occupancy grids with arbitrary user-specified resolutions. Optionally, to further mitigate the inherent noise in VFM predictions and tracking results, we introduce a Trilateral Radial Basis Function (TRBF), which jointly considers semantic, color, and spatial affinities to periodically smooth the Gaussian parameters. Extensive experiments on Occ3D-nuScenes [Occ3D] and the recently released high-resolution nuCraft [nuCraft] demonstrate that TT-Occ achieves better performance than existing self-supervised counterparts, which typically require hundreds of GPU training hours. Qualitative analysis further highlights the superiority of TT-Occ in terms of temporal consistency and open-vocabulary generalization.
@@
Figure 2: Overview of the proposed TT-OccCamera and $T$-OccLiDAR approaches. $T$-Occ performs test-time 3D occupancy estimation by directly integrating a suite of VFMs at runtime, avoiding any network training or fine-tuning. (a) Multi-view semantics are obtained using arbitrary open-vocabulary segmentation VFMs (OpenSeed, GroundingSAM, REX-Omni). (b) Geometry cues are extracted via depth/correspondence VFMs (VGGT, MapAnything). (c) Dynamic flow is estimated to track moving objects and prevent trailing artifacts while accumulating static structure over time. (d) The resulting features instantiate and refine time-aware 3D Gaussians, which can be voxelized at any resolutions for occupancy prediction. (e) $T$-Occ supports both LiDAR-based and camera-only variants and enables multi-resolution voxelization and semi-automatic annotation. Despite leveraging multiple VFMs, $T$-Occ remains highly efficient, delivering competitive occupancy performance across Occ3D-nuScenes and nuCraft.
@@ 2 Related Work
Self-Supervised Occupancy Prediction. Fully supervised occupancy methods predict voxel-level semantics using dense voxel grids [TPVFormer, PanoOcc, HTCL], depth priors [VoxFormer, Symphonize], or sparse representations [Octreeocc, SparseOcc, GaussianFormer]. Despite their effectiveness, these approaches rely heavily on costly large-scale 3D annotations. To mitigate this, recent methods explore self-supervised occupancy learning. SelfOcc [SelfOcc] leverages signed distance fields (SDF) and multi-view stereo embeddings to achieve temporally consistent occupancy from videos. OccNeRF [OccNeRF] utilizes photometric consistency and 2D foundation model supervision for semantic occupancy estimation in unbounded scenes. In open-world scenarios, POP3D [POP-3D] jointly trains class-agnostic occupancy grids and open-vocabulary semantics using unlabeled paired LiDAR and images, but suffers from sparsity and semantic ambiguity due to low-resolution CLIP [CLIP] features. VEON [VEON] introduces a vocabulary-enhanced occupancy framework trained with LiDAR supervision, leveraging CLIP features for open-vocabulary prediction and addressing depth ambiguities via enhance depth model (MiDaS [MiDaS], ZoeDepth [ZoeDepth]). GaussianOcc [GaussianOcc] uses Gaussian Splatting [3DGS] for cross-view optimization without pose annotations, while GaussianTR [GaussianTR] aligns rendered Gaussian features with pre-trained foundation models, enabling open-vocabulary occupancy prediction without explicit annotations. Despite these advances, existing methods either rely on extensive offline training or struggle with open-vocabulary settings and fixed resolutions. In contrast, our approach overcomes these limitations by enabling occupancy prediction through temporally coherent, training-free Gaussian optimization at test time.
@@
3D Reconstruction of Driving Scenes. Recent advances in dynamic scene modeling [vggt, szymanowicz2025flash3d, keetha2025mapanything] have achieved impressive photorealism and multi-view consistency. OmniRe [DriveStudio] performs real-time 3D reconstruction and simulation by building local canonical spaces for dynamic urban actors. Street Gaussians [Street] separates moving vehicles from static backgrounds, enabling efficient and high-quality rendering. DrivingGaussian [DrivingGaussian] incrementally reconstructs static scenes and dynamically integrates moving objects via Gaussian graphs for interactive editing. HUGS [HUGS] jointly optimizes geometry, appearance, semantics, and motion to achieve real-time view synthesis and 3D semantic reconstruction without explicit bounding box annotations. Autoregressive world modeling methods [OccWorld, OccSora] predict future occupancy using previously estimated 3D occupancies, facilitating temporal reasoning in dynamic environments. Our test-time approach fundamentally differs from these methods by eliminating dependencies on external priors and annotations (e.g., HD maps and GT bounding boxes). Instead, we focus solely on raw sensor inputs, optimizing Gaussian representations independently at each frame to directly infer the accurate geometry, rather than reconstructing photorealistic scenes or predicting future occupancy.
@@ 3.1 Lift Geometry and Semantics into Gaussians
Modality-Specific Initialization. In the LiDAR-based variant $\operatorname{TT-OccLiDAR}$, the sparse LiDAR points are directly initialized as 3D Gaussians, inheriting the precise spatial positions from real-world measurements. In contrast, the vision-centric variant $\operatorname{TT-OccCamera}$ reconstructs a 3D point cloud from 3D vision foundation models (3DVFMs). Specifically, we employ a pretrained 3DVFM such as VGGT [vggt] and MapAnything [keetha2025mapanything] to estimate dense depth maps from multi-view RGB inputs. When predicted depth maps suffer from inherent scale ambiguity due to the lack of metric supervision (e.g., in VGGT), we mitigate this issue by applying multi-view triangulation to keypoint correspondences in overlapping views. See Appendix for details.
VFM Semantics. To incorporate semantic information, we extract semantic maps from $M$ surrounding views by querying an open-vocabulary model [OpenSeeD, groundedsam, rexomni]. Our framework is agnostic to the specific VFM used and various segmentation and grounding models were evaluated, as illustrated in Fig. 2. These semantic maps are then lifted to 3D via a visibility-weighted projection:
@@
where $\mathbf{m}_{i}$ is a fused semantic probability vector for the 3D point. Specifically, $\bm{\mu_{i}}\in\mathbb{R}^{3}$ denotes the 3D coordinate of the $i$-th point. For each camera view $m\in\{1,\ldots,M\}$: $\mathrm{Proj}(\bm{\mu_{i}};\mathbf{K}_{m},\mathbf{E}_{m})$ projects the 3D point $\bm{\mu_{i}}$ onto the 2D image plane using the corresponding camera intrinsics $\mathbf{K}_{m}$ and extrinsics $\mathbf{E}_{m}$; $\mathcal{M}_{m}(\cdot)$ returns the semantic probability vector from the open-vocabulary segmentation map at the projected 2D location; $\mathbb{I}_{m}(\bm{\mu_{i}})$ is a binary visibility indicator, equal to 1 if the point $\bm{\mu_{i}}$ is visible in view $m$, and 0 otherwise. The use of foundation models enables compatibility with open-vocabulary semantic queries, allowing TT-Occ to flexibly adapt to user-specified class definitions at test time. For benchmark evaluation on nuScenes [nuScenes], we adopt the standard label space; however, our method inherently supports open-vocabulary settings without requiring retraining, in contrast to conventional self-supervised occupancy prediction approaches (e.g., [SelfOcc]) that depend on fixed decoder architectures and label sets.
Voxel-aware Simplifications. To accelerate Gaussian optimization and subsequent voxelization, we simplify the standard 3DGS [3DGS] by initializing the scale parameters with $\delta$ and constraining them using a sigmoid activation rather than an exponential function to prevent excessive growth. Additionally, we prune redundant Gaussians within the same voxel cell (size $\delta$) while merging their semantic probabilities.
@@ 3.2 Track Dynamic Gaussians
Reconstructing a driving scene faithfully can be challenging due to fast-moving objects (e.g., vehicles, pedestrians) that are often only partially observed. Without prior knowledge such as complete trajectories or bounding box annotations of moving instances used in [DrivingGaussian, DistillNeRF], optimizing 3D Gaussians online can often result in severe trailing artifacts. In the upper image of Fig. 3, fast-moving vehicles (blue voxels) produce noticeable trailing artifacts, while the lower image shows a clean reconstruction without such artifacts. To address this, we propose to track dynamic Gaussians while accumulating static ones across adjacent frames.
Modality-Specific Tracking. Both $\operatorname{TT-OccCamera}$ and $\operatorname{TT-OccLiDAR}$ share the same mechanism for static Gaussian inheritance, which enhances scene completeness by accumulating temporally consistent observations across frames. The key difference lies in how dynamic Gaussians are handled, particularly in their tracking strategy and purpose. For $\operatorname{TT-OccLiDAR}$, we track the Gaussian motion with learning-free Gaussian scene flow estimation, which allows us to relocate dynamic Gaussians accordingly. Our pipeline consists of the following steps: associating instances by projecting LiDAR points onto segmentation masks; denoising with DBSCAN [dbscan]; matching clusters across frames based on spatial proximity and shape similarity; and finally estimating 3D flow using the iterative closest point (ICP) algorithm [ICP]. See Appendix for implementation details.
For $\operatorname{TT-OccCamera}$, we estimate optical flow between adjacent frames of the same camera using RAFT [raft], and compute ego-motion flow based on inter-frame camera poses and per-pixel depth predicted by 3DVFMs (e.g., VGGT [vggt] or MapAnything [keetha2025mapanything]). Subtracting the ego-motion flow from the optical flow yields a residual dynamic flow, which reflects true object motion. Although this 2D dynamic flow could, in principle, guide the 3D motion of dynamic Gaussians, back-projecting it into 3D space tends to amplify noise from both RAFT and 3DVFMs, resulting in unstable Gaussian motion. To mitigate this, we adopt a compromise strategy by thresholding the dynamic flow magnitude to obtain a dynamic mask that identifies likely moving regions. The corresponding 3D Gaussians projected onto these regions are treated as dynamic and excluded from static accumulation in the next frame. While this approach does not allow accumulation of dynamic objects as in the LiDAR-based variant, it effectively reduces artifacts caused by noisy motion cues and temporal inconsistencies.
Table 1: Occupancy prediction performance on Occ3D-nuScenes [Occ3D]. Best results among self-supervised methods are highlighted in bold.
@@
$\operatorname{BEVFormer}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ECCV^{\prime}22)}}}$ [BEVFormer] C 26.88 37.83 17.87 40.44 42.43 7.36 23.88 21.81 20.98 22.38 30.70 55.35 36.0 28.06 20.04 17.69
@@
$\operatorname{CTF-Occ_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(NeurIPS^{\prime}23)}}}}$ [Occ3D] C 28.53 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 37.98 33.23 20.79 18.0
@@
$\operatorname{RenderOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ICRA^{\prime}24)}}}$ [RenderOcc] C 23.93 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 42.01 43.94 17.36 22.61
@@
$\operatorname{OccFlowNet}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(WACV^{\prime}25)}}}$ [OccFlowNet] C&L 26.14 27.50 26.00 34.00 32.00 20.40 25.90 18.60 20.20 26.00 28.70 62.00 37.80 39.50 29.00 26.80
@@
$\operatorname{SimpleOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(TIV^{\prime}24)}}}$ [SimpleOcc] C 7.99 0.67 1.18 3.21 7.63 1.02 0.26 1.80 0.26 1.07 2.81 40.44 18.30 17.01 13.42 10.84
@@
$\operatorname{OccNeRF}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(Arxiv^{\prime}24)}}}$ [OccNeRF] C 10.81 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 20.81 24.75 18.45 13.19
@@
$\operatorname{DistillNeRF}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(NeurIPS^{\prime}24)}}}$ [DistillNeRF] C 8.93 1.35 2.08 10.21 10.09 2.56 1.98 5.54 4.62 1.43 7.90 43.02 16.86 15.02 14.06 15.06

### v4

@@ Abstract
Self-supervised 3D occupancy prediction offers a promising solution for understanding complex driving scenes without requiring costly 3D annotations. However, training dense occupancy decoders to capture fine-grained geometry and semantics can demand hundreds of GPU hours, and once trained, such models struggle to adapt to varying voxel resolutions or novel object categories without extensive retraining. To overcome these limitations, we propose a practical and flexible test-time occupancy prediction framework termed TT-Occ. Our method incrementally constructs, optimizes, and voxelizes time-aware 3D Gaussians from raw sensor streams by integrating vision foundation models (VFMs) at runtime. The flexible representation of 3D Gaussians enables voxelization at arbitrary user-specified resolutions, while the strong generalization capability of VFMs supports accurate perception and open-vocabulary recognition without requiring any network training or fine-tuning. To validate the generality and effectiveness of our framework, we present two variants: a LiDAR-based version and a vision-centric version, and conduct extensive experiments on the Occ3D-nuScenes and nuCraft benchmarks under varying voxel resolutions. Experimental results show that TT-Occ significantly outperforms existing computationally expensive pretrained self-supervised counterparts. Code is available at https://github.com/Xian-Bei/TT-Occ.
@@ 1 Introduction
3D Occupancy prediction seeks to accurately identify regions within an environment that are occupied by objects of particular classes and those that remain free. This capability is crucial to enable collision-free trajectory planning and reliable navigation in autonomous driving systems [49, 44, 6, 7] and embodied agents [45, 40, 18, 31]. Existing solutions [13, 52, 27, 4, 43, 14, 35], however, heavily rely on dense 3D annotations obtained through labor-intensive manual labeling of dynamic driving scenes spanning up to 80 meters per frame. To mitigate this cost, recent studies have resorted to self-supervised alternatives [11, 50, 12, 39, 10, 16, 53, 3, 57]. While reducing labeling costs, they still require substantial computational overhead. For instance, training SelfOcc [12] on Occ3D-nuScenes [35] at a voxel resolution of 0.4m requires approximately 2 days on eight GPUs. Furthermore, once trained, adapting to finer resolution (e.g., 0.2m of nuCraft [58] dataset) or novel object classes (e.g., beyond the 17 predefined classes of Occ3D-nuScenes [35]) may necessitate extensive retraining.
@@
Recent vision foundation models (VFMs), however, change this landscape. Models such as VGGT [37] and MapAnything [19] provide reliable multiview geometry, and REX-Omni [17] enables open-vocabulary semantic reasoning, both directly at test time. As these cues become increasingly accessible without task-specific training, it is natural to revisit a long-standing assumption of the field: if such information no longer needs to be learned by a network, is training an occupancy model still necessary? To answer this question, we explore a test-time occupancy estimation method termed TT-Occ, which progressively constructs, optimizes and voxelizes time-aware 3D Gaussians from raw sensor streams by integrating VFMs. We introduce two variants, TT-OccCamera and TT-OccLiDAR, which differ in the sensor modality used to initialize the Gaussian primitives, respectively. Both variants eliminate costly pretraining and allow flexible adaptation to arbitrary voxel resolutions and user-specified semantic queries. Unlike previous NeRF- [47, 48] and 3DGS-based [56, 8] reconstruction methods that perform offline per-scene modeling assisted by external GT priors (e.g., HD maps or bounding boxes), TT-Occ generates occupancy representations in an online fashion, relying solely on raw sensor streams and generally trained VFMs to instantiate Gaussians capturing object geometry and semantics in unbounded outdoor scenes.
@@
Our framework follows a simple yet effective “lift-track-voxelize” symphony: (1) Lift: at each test time step, we first “lift” geometry and semantic information of surrounding views extracted via VFMs into time-aware 3D Gaussians on the fly. The generated Gaussians can also be splatted back onto the image plane through differentiable rasterization for parameter optimization [20]. (2) Track: next, we “track” dynamic Gaussians and accumulate static ones using estimated motion flow. This motion compensates for partial object visibility and prevents trailing artifacts while maintaining long-term temporal coherence. (3) Voxelize: at any given timestamp, the generated 3D Gaussians can be voxelized onto discrete occupancy grids with arbitrary user-specified resolutions. Optionally, to further mitigate the inherent noise in VFM predictions and tracking results, we introduce a Trilateral Radial Basis Function (TRBF), which jointly considers semantic, color, and spatial affinities to periodically smooth the Gaussian parameters.
Extensive experiments on Occ3D-nuScenes [35] and the recently released high-resolution nuCraft [58] demonstrate that TT-Occ achieves better performance than existing self-supervised counterparts, which typically require hundreds of GPU training hours. Qualitative analysis further highlights the superiority of TT-Occ in terms of temporal consistency and open-vocabulary generalization.
@@
Figure 2: Overview of the proposed TT-OccCamera and $T$-OccLiDAR approaches. $T$-Occ performs test-time 3D occupancy estimation by directly integrating a suite of VFMs at runtime, avoiding any network training or fine-tuning. (a) Multi-view semantics are obtained using arbitrary open-vocabulary segmentation VFMs (e.g., OpenSeed, GroundingSAM, REX-Omni). (b) Geometry cues are extracted via depth/correspondence VFMs (e.g., VGGT, MapAnything). (c) Dynamic flow is estimated to track moving objects and prevent trailing artifacts while accumulating static structure over time. (d) The resulting features instantiate and refine time-aware 3D Gaussians, which can be voxelized at any resolutions for occupancy prediction. (e) $T$-Occ supports both LiDAR-based and camera-only variants and enables multi-resolution voxelization and semi-automatic annotation. Despite leveraging multiple VFMs, $T$-Occ remains highly efficient, delivering competitive occupancy performance across Occ3D-nuScenes and nuCraft.
@@ 2 Related Work
Occupancy Prediction. Fully supervised occupancy methods predict voxel-level semantics using dense voxel grids [13, 42, 21], depth priors [22, 15], or sparse representations [25, 33, 14]. Despite their effectiveness, these approaches rely heavily on costly large-scale 3D annotations. To mitigate this, recent methods explore self-supervised occupancy learning. SelfOcc [12] leverages signed distance fields (SDF) and multi-view stereo embeddings to achieve temporally consistent occupancy from videos. OccNeRF [50] utilizes photometric consistency and 2D foundation model supervision for semantic occupancy estimation in unbounded scenes. In open-world scenarios, POP3D [36] jointly trains class-agnostic occupancy grids and open-vocabulary semantics using unlabeled paired LiDAR and images, but suffers from sparsity and semantic ambiguity due to low-resolution CLIP [28] features. VEON [53] introduces a vocabulary-enhanced occupancy framework trained with LiDAR supervision, leveraging CLIP features for open-vocabulary prediction and addressing depth ambiguities via enhance depth model (MiDaS [29], ZoeDepth [2]). GaussianOcc [10] uses Gaussian Splatting [20] for cross-view optimization without pose annotations, while GaussianTR [16] aligns rendered Gaussian features with pre-trained foundation models, enabling open-vocabulary occupancy prediction without explicit annotations. Despite these advances, existing methods either rely on extensive offline training or struggle with open-vocabulary settings and fixed resolutions. In contrast, our approach overcomes these limitations by enabling occupancy prediction through temporally coherent, training-free Gaussian optimization at test time.
@@
3D Reconstruction of Driving Scenes. Recent advances in dynamic scene modeling [37, 32, 19] have achieved impressive photorealism and multi-view consistency. OmniRe [8] performs real-time 3D reconstruction and simulation by building local canonical spaces for dynamic urban actors. Street Gaussians [46] separates moving vehicles from static backgrounds, enabling efficient and high-quality rendering. DrivingGaussian [56] incrementally reconstructs static scenes and dynamically integrates moving objects via Gaussian graphs for interactive editing. HUGS [55] jointly optimizes geometry, appearance, semantics, and motion to achieve real-time view synthesis and 3D semantic reconstruction without explicit bounding box annotations. Autoregressive world modeling methods [54, 38] predict future occupancy using previously estimated 3D occupancies, facilitating temporal reasoning in dynamic environments. Our test-time approach fundamentally differs from these methods by eliminating dependencies on external priors and annotations (e.g., HD maps and GT bounding boxes). Instead, we focus solely on raw sensor inputs, optimizing Gaussian representations independently at each frame to directly infer the accurate geometry, rather than reconstructing photorealistic scenes or predicting future occupancy.
@@ 3.1 Lift Geometry and Semantics into Gaussians
Modality-Specific Initialization. In the LiDAR-based variant $\operatorname{TT-OccLiDAR}$, the sparse LiDAR points are directly initialized as 3D Gaussians, inheriting the precise spatial positions from real-world measurements. In contrast, the vision-centric variant $\operatorname{TT-OccCamera}$ reconstructs a 3D point cloud from 3D vision foundation models (3DVFMs). Specifically, we employ a pretrained 3DVFM such as VGGT [37] and MapAnything [19] to estimate dense depth maps from multi-view RGB inputs. When predicted depth maps suffer from inherent scale ambiguity due to the lack of metric supervision (e.g., in VGGT), we mitigate this issue by applying multi-view triangulation to keypoint correspondences in overlapping views. See supplementary materials for details.
VFM Semantics. To incorporate semantic information, we extract semantic maps from $M$ surrounding views by querying an open-vocabulary model [51, 30, 17]. Our framework is agnostic to the specific VFM used and various segmentation and grounding models were evaluated, as illustrated in Fig. 2. These semantic maps are then lifted to 3D via a visibility-weighted projection:
@@
where $\mathbf{m}_{i}$ is a fused semantic probability vector for the 3D point. Specifically, $\bm{\mu_{i}}\in\mathbb{R}^{3}$ denotes the 3D coordinate of the $i$-th point. For each camera view $m\in\{1,\ldots,M\}$: $\mathrm{Proj}(\bm{\mu_{i}};\mathbf{K}_{m},\mathbf{E}_{m})$ projects the 3D point $\bm{\mu_{i}}$ onto the 2D image plane using the corresponding camera intrinsics $\mathbf{K}_{m}$ and extrinsics $\mathbf{E}_{m}$; $\mathcal{M}_{m}(\cdot)$ returns the semantic probability vector from the open-vocabulary segmentation map at the projected 2D location; $\mathbb{I}_{m}(\bm{\mu_{i}})$ is a binary visibility indicator, equal to 1 if the point $\bm{\mu_{i}}$ is visible in view $m$, and 0 otherwise. The use of foundation models enables compatibility with open-vocabulary semantic queries, allowing TT-Occ to flexibly adapt to user-specified class definitions at test time. For benchmark evaluation on nuScenes [5], we adopt the standard label space; however, our method inherently supports open-vocabulary settings without requiring retraining, in contrast to conventional self-supervised occupancy prediction approaches (e.g., [12]) that depend on fixed decoder architectures and label sets.
Voxel-aware Simplifications. To accelerate Gaussian optimization and subsequent voxelization, we simplify the standard 3DGS [20] by initializing the scale parameters with $\delta$ and constraining them using a sigmoid activation rather than an exponential function to prevent excessive growth. Additionally, we prune redundant Gaussians within the same voxel cell (size $\delta$) while merging their semantic probabilities.
@@ 3.2 Track Dynamic Gaussians
Reconstructing a driving scene faithfully can be challenging due to fast-moving objects (e.g., vehicles, pedestrians) that are often only partially observed. Without prior knowledge such as complete trajectories or bounding box annotations of moving instances used in [56, 39], optimizing 3D Gaussians online can often result in severe trailing artifacts. To address this, we propose to track dynamic Gaussians while accumulating static ones across adjacent frames. In the upper image of Fig. 3, fast-moving vehicles (blue voxels) produce noticeable trailing artifacts, while the lower image shows a clean reconstruction after the Gaussian tracking.
Modality-Specific Tracking. Both $\operatorname{TT-OccCamera}$ and $\operatorname{TT-OccLiDAR}$ share the same mechanism for static Gaussian inheritance, which enhances scene completeness by accumulating temporally consistent observations across frames. The key difference lies in how dynamic Gaussians are handled, particularly in their tracking strategy and purpose. For $\operatorname{TT-OccLiDAR}$, we track the Gaussian motion with learning-free Gaussian scene flow estimation, which allows us to relocate dynamic Gaussians accordingly. Our pipeline consists of the following steps: associating instances by projecting LiDAR points onto segmentation masks; denoising with DBSCAN [9]; matching clusters across frames based on spatial proximity and shape similarity; and finally estimating 3D flow using the iterative closest point (ICP) algorithm [1]. For $\operatorname{TT-OccCamera}$, we estimate optical flow between adjacent frames of the same camera using RAFT [34], and compute ego-motion flow based on inter-frame camera poses and per-pixel depth predicted by 3DVFMs (e.g., VGGT [37] or MapAnything [19]). Subtracting the ego-motion flow from the optical flow yields a residual dynamic flow, which reflects true object motion. Although this 2D dynamic flow could, in principle, guide the 3D motion of dynamic Gaussians, back-projecting it into 3D space tends to amplify noise from both RAFT and 3DVFMs, resulting in unstable Gaussian motion. To mitigate this, we adopt a compromise strategy by thresholding the dynamic flow magnitude to obtain a dynamic mask that identifies likely moving regions. The corresponding 3D Gaussians projected onto these regions are treated as dynamic and excluded from static accumulation in the next frame. While this approach does not allow accumulation of dynamic objects as in the LiDAR-based variant, it effectively reduces artifacts caused by noisy motion cues and temporal inconsistencies. See supplementary materials for implementation details.
Table 1: Occupancy prediction performance on Occ3D-nuScenes [35]. Best results among self-supervised methods are highlighted in bold.
@@
$\operatorname{BEVFormer}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ECCV^{\prime}22)}}}$ [23] C 26.88 37.83 17.87 40.44 42.43 7.36 23.88 21.81 20.98 22.38 30.70 55.35 36.0 28.06 20.04 17.69
@@
$\operatorname{CTF-Occ_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(NeurIPS^{\prime}23)}}}}$ [35] C 28.53 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 37.98 33.23 20.79 18.0
@@
$\operatorname{RenderOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(ICRA^{\prime}24)}}}$ [27] C 23.93 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 42.01 43.94 17.36 22.61
@@
$\operatorname{OccFlowNet}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(WACV^{\prime}25)}}}$ [4] C&L 26.14 27.50 26.00 34.00 32.00 20.40 25.90 18.60 20.20 26.00 28.70 62.00 37.80 39.50 29.00 26.80
@@
$\operatorname{SimpleOcc}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(TIV^{\prime}24)}}}$ [11] C 7.99 0.67 1.18 3.21 7.63 1.02 0.26 1.80 0.26 1.07 2.81 40.44 18.30 17.01 13.42 10.84
@@
$\operatorname{OccNeRF}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(Arxiv^{\prime}24)}}}$ [50] C 10.81 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 20.81 24.75 18.45 13.19
@@
$\operatorname{DistillNeRF}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\small\operatorname{(NeurIPS^{\prime}24)}}}$ [39] C 8.93 1.35 2.08 10.21 10.09 2.56 1.98 5.54 4.62 1.43 7.90 43.02 16.86 15.02 14.06 15.06